# Poisoning Prevention in Federated Learning and Differential Privacy via Stateful Proofs of Execution

### v2

@@ IX-D Time Overhead
As $\sf{\mathcal{V}rf}$ operates on a powerful back-end, its runtime within the $\sf{SLAPP}$ protocol is negligible compared to the execution time on $\sf{\mathcal{P}rv}$. Thus, we focus on measuring $\sf{\mathcal{P}rv}$’s execution time (i.e., the runtime time of executing Phases 3, 4, and 5 in Section V). Results are illustrated in Figure 5.
As expected, $\sf{SLAPP}_{SK}$ has the fastest runtime: less than half a second. On the other hand, $\sf{SLAPP}_{PQ}$ utilizes expensive cryptography and thus incurs the longest execution time of $\approx 100$ seconds. $\sf{SLAPP}_{PK}$ positions between these two, taking around 2.5 seconds.
@@
Since $\sf{SLAPP}$ requires $\sf{\mathcal{P}rv}$ to compute a snapshot of $PMEM$ as part of proof generation, we conducted an experiment to assess the impact of $PMEM$ size on $\sf{\mathcal{P}rv}$’s execution time. In this experiment, we varied the $PMEM$ size from 4KB to 32KB. As depicted in Figure 6, the time to execute this snapshot contributes significantly to the overall execution time when $\sf{SLAPP}_{SK}$ is used. It also shows a linear relationship between execution time and $PMEM$ size for $\sf{SLAPP}_{SK}$. However, for $\sf{SLAPP}_{PK}$ and $\sf{SLAPP}_{PQ}$, this effect is negligible; its runtime has minimal impact on the overall execution time regardless of $PMEM$ size.
Also, from Figures 5 and 6, we can conclude that all $\sf{SLAPP}$ variants incur a negligible execution time overhead, i.e., $<1\%$, compared to their baseline counterparts.
@@ IX-E Case Study 1: Local Differential Privacy
In the first case study, we envision the integration of $\sf{SLAPP}$ within a smart city/grid system. The service provider aims to periodically collect energy consumption data from all smart meters located in individual households to calculate electric bills, forecast load, etc. Previous studies [35, 38] have demonstrated privacy risks by exposing raw energy to the service provider, e.g., with access to such data, the provider could potentially infer users’ habits and behaviors.
To address this privacy concern, one can employ LDP-DC from Algorithm 1. The service provider is also motivated to employ a poisoning-free version, ${}^{+}$ as presented in Section VIII-A, to prevent poisoning by detecting malicious outputs from edge devices.
TABLE II: Performance of FL/LDP with $\sf{SLAPP}_{SK}$
Protocol Task Phase Runtime Non-Secure Secure
Mem. Usage Mem. Usage
${}^{+}$ Collect noisy Setup 0.50s 1KB 1.48KB
energy readings Collect 2.30s 1.27KB 1.48KB
${}^{+}$ Train a classifier for detecting anomalous energy usage patterns Setup 0.50s 2.57KB 1.48KB
Collect 1.98s 2.7KB 1.48KB
Train 21.84s 3.97KB 1.48KB
@@
We build a prototype of a smart meter ($\sf{\mathcal{P}rv}$) based on a SAML11 Xplained Pro development board connecting to a PZEM-004T energy meter module. $\sf{\mathcal{P}rv}$’s Non-Secure World consists of LDP-DC software (Algorithm 1) and a driver responsible for retrieving energy data from the PZEM-004T module. Meanwhile, the Secure World contains $\sf{SLAPP}_{SK}$’s TCB. The service provider ($\sf{\mathcal{V}rf}$) runs on a commodity desktop and communicates with $\sf{\mathcal{P}rv}$ over serial communication.
@@
During normal (benign) operation, $\sf{\mathcal{V}rf}$ and $\sf{\mathcal{P}rv}$ execute an instance of ${}^{+}$ protocol. This instance begins with the Setup phase, which requires $\sf{PoSX}$ of Init-State function to initialize a $\sf{\mathcal{P}rv}$-local state ($PRR$ mapping) to zeroes. Since $\sf{\mathcal{P}rv}$ faithfully executes this phase, it completes with $\top$. Upon obtaining the successful output, $\sf{\mathcal{V}rf}$ proceeds to the Collect phase, which aims to collect authentic noisy energy data from $\sf{\mathcal{P}rv}$. To achieve this, $\sf{\mathcal{V}rf}$ crafts a $\sf{PoSX}$ request by configuring a function to prove execution to LDP-DC and properly selecting LDP input parameters (i.e., $f$, $p$ and $q$) to meet the $\epsilon$-LDP requirement, and sends an authenticated $\sf{PoSX}$ request to $\sf{\mathcal{P}rv}$. Adhering to the protocol, $\sf{\mathcal{P}rv}$ executes LDP-DC correctly (with $\sf{\mathcal{V}rf}$-defined inputs) and returns authentic noisy energy data to $\sf{\mathcal{V}rf}$. As a result, this phase completes with $\top$ ensuring $\sf{\mathcal{V}rf}$ that the received energy data is not poisoned. End-to-end runtime and memory usage during this operation are shown in Table II.
Next, we launch the following attacks to ${}^{+}$:
@@
${}^{+}$ detects all aforementioned attacks. Specifically, ${\sf{\mathcal{A}dv}}_{1}$ and ${\sf{\mathcal{A}dv}}_{2}$ are detected at the end of Setup and Collect phases, respectively, due to the software binary mismatch. ${\sf{\mathcal{A}dv}}_{3}$ is caught in Collect since $\sf{SLAPP}$’s RoT noticed tampered state values (from CheckState) and thus refused to generate a valid proof. Similarly, ${\sf{\mathcal{A}dv}}_{4}$ is detected in Collect phase since the proof $\sigma$ does not reflect the tampered output.
@@ IX-F Case Study 2: Federated Learning
Here the service provider aims to train an ML model for identifying anomalous usage patterns in a smart grid system. To achieve this while ensuring user privacy, it employs FL based on a neural network, as described in [27]. Each smart meter independently trains a local neural network model with 8 hidden units on its local data. These local models are later transmitted to and aggregated by the service provider. ${}^{+}$ is employed to prevent poisoning.
The setup resembles the previous case study, where a smart meter device ($\sf{\mathcal{P}rv}$) operates on a SAML11 Xplained Pro board. It performs energy readings by interfacing with a PZEM-004T device and connects to $\sf{\mathcal{V}rf}$ via serial communication.
@@
During benign operation, $\sf{\mathcal{P}rv}$ and $\sf{\mathcal{V}rf}$ initiate ${}^{+}$ protocol by running Setup phase, which clears all sensor readings on $\sf{\mathcal{P}rv}$. Next, $\sf{\mathcal{V}rf}$ periodically triggers Collect phase to record an energy reading into a $\sf{\mathcal{P}rv}$-local training dataset. Once $\sf{\mathcal{P}rv}$ records enough energy readings, $\sf{\mathcal{V}rf}$ triggers the Train phase by transmitting a $\sf{PoSX}$ request to $\sf{\mathcal{P}rv}$. This request specifies initial weights to be trained on, a learning rate, and the number of epochs to be 100. $\sf{\mathcal{P}rv}$ then performs the local training. Upon completion, $\sf{\mathcal{P}rv}$ sends back the updated weights along with the proof of training to $\sf{\mathcal{V}rf}$. Since $\sf{\mathcal{V}rf}$ and $\sf{\mathcal{P}rv}$ adhere to ${}^{+}$, the protocol yields $\top$, ensuring to $\sf{\mathcal{V}rf}$ that the received model was correctly trained on authentic energy data and can thus be safely aggregated onto the global model. Runtime breakdown and memory usage are shown in Table II.
@@
We simulate two types of adversaries: ${\sf{\mathcal{A}dv}}_{D}$ and ${\sf{\mathcal{A}dv}}_{M}$. ${\sf{\mathcal{A}dv}}_{D}$ performs data poisoning attacks to ${}^{+}$ by compromising $\sf{\mathcal{P}rv}$-local training dataset before Train is executed. Meanwhile, ${\sf{\mathcal{A}dv}}_{M}$ simulates model poisoning by manipulating the locally trained model parameters before the model arrives at $\sf{\mathcal{V}rf}$. ${}^{+}$ detects ${\sf{\mathcal{A}dv}}_{D}$ during the Train phase because $\sf{SLAPP}$’s RoT identifies state $\sf{\mathcal{S}}$ tampering ($\sf{\mathcal{S}}$ is the training dataset in this case). ${\sf{\mathcal{A}dv}}_{M}$ is caught at the end of the Train as the proof does not match the tampered output.
@@ X Related Work
Poisoning prevention LDP and FL. Aside from data-driven  [57, 23, 6]) and algorithmic [50, 32, 55, 33] poisoning prevention approaches discussed in Section I, a few recent schemes propose system security approaches [11, 48, 2]. To the best of our knowledge, none of these schemes provide $\sf{PoSX}$-equivalent guarantee of provable integrity all the way from data acquisition until its de facto usage as an aggregated statistical result or global ML model. They also do not address the issues of input validation or state preservation for arbitrary functions $\sf{\mathcal{F}}$ to be execute on $\sf{\mathcal{P}rv}$.
@@
Private Data Collection on Edge Devices. Complementary to privacy mechanisms focusing on hiding private data from back-ends (e.g., LDP/FL), recent work has delved into assuring that private data is secure against compromised sensing devices “from its birth”, i.e., from the moment when it is digitized. VERSA [43] was proposed as a HW/SW architecture to guarantee that only the correct execution of expected and explicitly authorized software can access and manipulate sensing interfaces. As a consequence, it blocks malware/modified software from accessing sensitive sensed quantities by default. Following this notion, Sensing And Actuation As A Privilege (SA4P) [16] realizes this concept using ARM TrustZone.

### v3

@@ IX-D Time Overhead
As $\sf{\mathcal{V}rf}$ operates on a powerful back-end, its runtime within the $\sf{SLAPP}$ protocol is negligible compared to the execution time on $\sf{\mathcal{P}rv}$. Thus, we focus on measuring $\sf{\mathcal{P}rv}$’s execution time (i.e., the runtime time of executing Phases 3, 4, and 5 in Section V). Results are illustrated in Fig. 3(b).
[…]
@@
Since $\sf{SLAPP}$ requires $\sf{\mathcal{P}rv}$ to compute a snapshot of $PMEM$ as part of proof generation, we conducted an experiment to assess the impact of $PMEM$ size on $\sf{\mathcal{P}rv}$’s execution time. In this experiment, we varied the $PMEM$ size from 32KB to 256KB. As depicted in Fig. 4, the time to execute this snapshot contributes significantly to the overall execution time when $\sf{SLAPP}_{SK}$ is used. It also shows a linear relationship between execution time and $PMEM$ size for $\sf{SLAPP}_{SK}$ and $\sf{SLAPP}_{PK}$. However, for $\sf{SLAPP}_{PQ}$, this effect is negligible; its runtime has minimal impact on the overall execution time regardless of $PMEM$ size.
From Figs. 3(b) and 4, we can conclude that all $\sf{SLAPP}$ variants incur a negligible execution time overhead, i.e., $<1\%$, compared to their baseline counterparts. We next conduct end-to-end evaluation of $\sf{SLAPP}$ through two case studies.
@@ IX-E Case Study 1: Local Differential Privacy
Description. In the first case study, we envision the integration of $\sf{SLAPP}$ within a smart city/grid system. The service provider aims to periodically collect energy consumption data from all smart meters located in individual households to calculate electric bills, forecast load, etc. Previous studies [36, 38] have demonstrated privacy risks by exposing raw energy to the service provider, e.g., with access to such data, the provider could potentially infer users’ habits and behavior. LDP-DC (from Algorithm 1) can be employed to address this concern. The service provider is also motivated to use a poisoning-free version, LDP-DC+ as presented in Section VIII-A, to prevent poisoning attacks from potentially malicious edge devices.
TABLE II: End-to-end results for $\sf{Baseline_{SK}}$ (shown in parentheses) and $\sf{SLAPP}_{SK}$. Phases 1-5 are as defined in Section V-B.
Case Study Protocol Step Average Execution Time (ms) RAM (KB) FLASH (KB)
Phase 1 Phase 2 Phase 3 Phase 4 Phase 5 Total S NS S NS
(1) Collect noisy energy readings LDP-DC+ Setup 8 660 7 1 11 687(678) 4.7(4.9) 2.7(1.2) 17.4(23.2) 8.5(7.4)
Collect 6 660 9 1 8 684(677) 4.7(4.9) 2.7(1.2)
(2) Train a model for load forecasting FL-DC+ Setup 13 652 8 1 8 682(678) 4.7(5.4) 3.3(1.5) 17.4(71.3) 57.2(7.5)
Collect 12 654 8 1 7 682(682) 4.7(5.4) 3.3(1.5)
Train 16 651 3,067 2 20 3,756(3,759) 4.7(32.0) 32.0(1.5)
@@
We build a prototype of a smart meter ($\sf{\mathcal{P}rv}$) based on NUCLEO-L552ZE-Q connecting to a PZEM-004T energy meter hardware module. $\sf{\mathcal{P}rv}$’s Non-Secure World consists of LDP-DC software (Algorithm 1) and a driver responsible for retrieving energy data from the PZEM-004T hardware module. Meanwhile, the Secure World contains $\sf{SLAPP}_{SK}$’s TCB. The service provider ($\sf{\mathcal{V}rf}$) runs on a commodity desktop and communicates with $\sf{\mathcal{P}rv}$ over serial communication.
@@
During normal (benign) operation, $\sf{\mathcal{V}rf}$ and $\sf{\mathcal{P}rv}$ execute an instance of LDP-DC+ protocol. This instance begins with the Setup step, which requires $\sf{PoSX}$ of Init-State function to initialize a $\sf{\mathcal{P}rv}$-local state ($PRR$ mapping) to zeroes. $\sf{\mathcal{V}rf}$ verified that $\sf{\mathcal{P}rv}$ faithfully executes this step, completing with $\top$. Upon obtaining the successful output, $\sf{\mathcal{V}rf}$ proceeds to the Collect, which aims to collect authentic noisy energy data from $\sf{\mathcal{P}rv}$. To achieve this, $\sf{\mathcal{V}rf}$ makes a $\sf{PoSX}$ request by configuring a function to prove execution to LDP-DC and properly selecting LDP input parameters (i.e., $f$, $p$ and $q$) to meet the $\epsilon$-LDP requirement, and sends an authenticated $\sf{PoSX}$ request to $\sf{\mathcal{P}rv}$. To successfully respond to this $\sf{PoSX}$ request, $\sf{\mathcal{P}rv}$ must execute LDP-DC correctly (with $\sf{\mathcal{V}rf}$-defined inputs) and return the result (authentic noisy energy data) to $\sf{\mathcal{V}rf}$. If this step completes with $\top$, it ensures $\sf{\mathcal{V}rf}$ that the received energy data is not poisoned.
End-to-end Evaluation. Results for this case study are presented in Table II. $\sf{SLAPP}_{SK}$ takes around 17.4KB and 8.5KB of Secure and Non-Secure FLASH, respectively. $\sf{Baseline_{SK}}$ would require placing the code implementing the LDP logics into Secure FLASH, enlarging its TCB by 5.8KB or 33.3%. Note that this number would further increase with complexity of functionality or if multiple different sensing functions need to be implemented by $\sf{\mathcal{P}rv}$. This result substantiates $\sf{SLAPP}$’s design rationale: being application-agnostic significantly reduces the TCB size and provides flexibility. Regarding RAM usage, $\sf{SLAPP}_{SK}$ requires 4.7KB of RAM to execute the Setup and Collect steps, which is 0.2KB less than $\sf{Baseline_{SK}}$, leading to 4.3% reduction in Secure RAM usage.
For both steps, the end-to-end execution time (from $\sf{\mathcal{V}rf}$ generating a request to $\sf{\mathcal{V}rf}$ verifying the response) is approximately 0.69 seconds, corresponding to 1% increase over $\sf{Baseline_{SK}}$. We also reported the execution time breakdown for all phases in Table II. Recall Section V-B for the definition of each phase. As Phase 2 requires a hash computation over the Non-Secure FLASH of 256KB, it dominates the end-to-end runtime, accounting for $\approx$96of the overall runtime. Phase 4 is the fastest, as it consists of only lightweight operations, i.e., setting flags and computing HMAC on a short message, without requiring any cross-world switching or network communication. Finally, the time taken by $\sf{\mathcal{V}rf}$ (Phases 1 and 5) contributes with $<$ 3% of the overall runtime.
We next consider a task of using LDP-DC+ to collect a variable number of noisy energy readings from 1 to 20. This emulates the case of continuous data collection to be performed over a longer period of time. For instance, $\sf{\mathcal{V}rf}$ requests 1 reading every 3 minutes, resulting in 20 readings over an hour. The results are shown in Fig. 5(a). As this task requires invoking Setup once before Collect can be repeated as many times as needed, the time for Setup remains constant, irrespective of the number of subsequent collections. Conversely, the runtime overhead for completing all Collect steps increases linearly with the number of readings collected. We do not observe a significant runtime overhead of $\sf{SLAPP}_{SK}$ compared to $\sf{Baseline_{SK}}$ in this case.
Attack Simulation. We launch the following attacks to LDP-DC+ in this case study:
@@
LDP-DC+ detects all aforementioned attacks. Specifically, ${\sf{\mathcal{A}dv}}_{1}$ and ${\sf{\mathcal{A}dv}}_{2}$ are detected at the end of Setup and Collect phases, respectively. ${\sf{\mathcal{A}dv}}_{3}$ is caught in Collect since $\sf{SLAPP}$’s RoT noticed tampered state values (from CheckState) and thus refused to generate a valid proof. Similarly, ${\sf{\mathcal{A}dv}}_{4}$ is detected in Collect phase since the proof $\sigma$ does not reflect the tampered output received by $\sf{\mathcal{V}rf}$.
(a) Case study 1: LDP-DC+
(b) Case study 2: FL-DC+
Figure 5: Sum of the runtimes of multiple data collection rounds performed over a longer period with multiple readings.
@@ IX-F Case Study 2: Federated Learning
Description. Here $\sf{\mathcal{V}rf}$ aims to train a ML model for one-hour-ahead load forecasting [54]. To achieve this while ensuring user privacy, it employs FL based on LSTM [54]. Each $\sf{\mathcal{P}rv}$ independently trains a lightweight LSTM model with one layer of 8 neurons on hourly energy readings. This model is later transmitted to and aggregated by $\sf{\mathcal{V}rf}$. FL-DC+ is employed to prevent poisoning.
@@
$\sf{\mathcal{P}rv}$ and $\sf{\mathcal{V}rf}$ initiate FL-DC+ protocol by running Setup step, which clears all energy readings on $\sf{\mathcal{P}rv}$. Next, $\sf{\mathcal{V}rf}$ periodically requests Collect to record an energy reading into a $\sf{\mathcal{P}rv}$-local training dataset. Once $\sf{\mathcal{P}rv}$ records enough energy readings, $\sf{\mathcal{V}rf}$ triggers the Train phase by transmitting a $\sf{PoSX}$ request to $\sf{\mathcal{P}rv}$. This request specifies the initial weights to be trained on, the learning rate (0.01), and the number of epochs (5). $\sf{\mathcal{P}rv}$ then (provably) performs the local training. Upon completion, $\sf{\mathcal{P}rv}$ sends back the updated weights along with the proof of training to $\sf{\mathcal{V}rf}$. Since $\sf{\mathcal{V}rf}$ and $\sf{\mathcal{P}rv}$ adhere to FL-DC+, the protocol yields $\top$, ensuring to $\sf{\mathcal{V}rf}$ that the received model was correctly trained on authentic energy data and thus can be securely aggregated onto the global model.
End-to-end Evaluation. The results of this case study are shown in Table II. Similar to the previous case study, by making the TCB application-agnostic, $\sf{SLAPP}_{SK}$ can reduce the size of Secure FLASH by a significant amount (76%). $\sf{SLAPP}_{SK}$ requires the same amount of Secure RAM regardless of the steps. In contrast, $\sf{Baseline_{SK}}$ incurs 0.7KB of Secure RAM for the Setup and Collect steps, while the Train step, which involves training an LSTM model, requires a more substantial amount, 32KB, of Secure RAM – around 5x of $\sf{SLAPP}_{SK}$. These results further emphasize $\sf{SLAPP}$’s greater benefits especially when $\sf{\mathcal{F}}$ is more resource-intensive, as in the case of LSTM training.
In terms of execution time, the results for the Setup and Collect steps are similar to the previous case study, with Phase 2 being the most time-consuming. However, the end-to-end execution time for Train is dominated by Phase 3, which involves executing $\sf{\mathcal{F}}$ (LSTM training) in the Non-Secure World. This step takes around 3 seconds, contributing $\approx$81to the end-to-end runtime.
Finally, we consider a task of applying FL-DC+ to collaboratively train an LSTM model on hourly energy readings collected over different numbers of days (from 1 to 20 days). The sums of total runtimes for the entire periods are shown in Fig. 5(b). Since Setup is performed once, its overall runtime is fixed to 0.69s. As the energy sensor is read 24 times a day, the time sum for all Collect steps is linear with the number of days and dominates the overall runtime. The Train step is executed at the end of this task on all previously collected readings, resulting in a linear runtime sum ranging from $2s$ (over 2 days) to $12s$ (over 20 days). With most steps exhibiting a linear runtime, the total runtime of this task also becomes linear with the number of days.
@@
Attack Simulation. We simulate two types of adversaries: ${\sf{\mathcal{A}dv}}_{D}$ and ${\sf{\mathcal{A}dv}}_{M}$. ${\sf{\mathcal{A}dv}}_{D}$ performs data poisoning attacks to FL-DC+ by compromising $\sf{\mathcal{P}rv}$-local training dataset before Train is executed. Meanwhile, ${\sf{\mathcal{A}dv}}_{M}$ simulates model poisoning by manipulating the locally trained model parameters before the model arrives at $\sf{\mathcal{V}rf}$. FL-DC+ detects ${\sf{\mathcal{A}dv}}_{D}$ during the Train phase because $\sf{SLAPP}$’s RoT identifies state $\sf{\mathcal{S}}$ tampering ($\sf{\mathcal{S}}$ is the training dataset in this case). ${\sf{\mathcal{A}dv}}_{M}$ is caught at the end of the Train as the proof does not match the tampered output.
@@ X Related Work
TABLE III: Comparison with current defenses ( ,  and  indicates the degree of support/assumption/impact)
Feature ($\rightarrow$) Approach ($\downarrow$) $\sf{\mathcal{A}dv}$ Data Application $\sf{\mathcal{P}rv}$
Adaptive
Robustness
Byzantine
Assumption
Distribution
Assumption
Impact on
Agg. Utility
Support
DP/LDP
Support
FL/ML
Require
RoT
App-indep.
RoT
Data-driven Sniper [5] N/A
Normalize. [6] N/A
ERR+LFR [22] N/A
Clustering [33] N/A
Algorithmic Multi-Krum [3] N/A
RoLR [23] N/A
TRIM [28] N/A
FL-WBC [52]
System-level CrowdGuard [49]
$\sf{SLAPP}$ (this work)
Poisoning prevention in LDP and FL. We divide existing approaches to preventing poisoning attacks into 3 categories: data-driven, algorithmic and system-level. Data-driven approaches detect poisoning attacks based solely on the collected data without modifying the underlying collection scheme, e.g., by applying normalization [6] and analyzing distances between or error rates from data [5, 33, 22]. Meanwhile, algorithmic approaches modify the data collection algorithm to make it resilient against poisoning attacks, e.g., by incorporating with Byzantine fault-tolerant techniques [3, 24] or using modified loss functions [23, 28, 52]. Finally, system-level approaches [49] leverage $\sf{\mathcal{P}rv}$’s security architecture to mitigate poisoning attacks. $\sf{SLAPP}$ falls into the last category.
Based on these categories, we qualitatively compare $\sf{SLAPP}$ with current defenses in Table III. It shows that $\sf{SLAPP}$ is the only approach that offers strong robustness against adaptive adversaries without making assumptions about Byzantine $\sf{\mathcal{A}dv}$ (i.e., $\sf{SLAPP}$ allows $\sf{\mathcal{A}dv}$ to corrupt any number of $\sf{\mathcal{P}rv}$-s) or data distributions (i.e., $\sf{SLAPP}$ allows each $\sf{\mathcal{P}rv}$ to have independent data distributions). As $\sf{SLAPP}$ can deterministically discern poisoned data from the benign, it results in no utility loss on aggregate data. It also supports both LDP and FL-based data collection schemes. As a system-level approach, $\sf{SLAPP}$ requires $\sf{\mathcal{P}rv}$ to have an RoT, which is becoming more common in modern IoT devices. Moreover, $\sf{SLAPP}$’s RoT hosts only a small TCB agnostic to the data collection scheme (FL or LDP); this enables a one-and-done process for validating its correctness (e.g., through formal verification, which is an interesting avenue for future work). Besides, to the best of our knowledge, none of existing work provides $\sf{PoSX}$-equivalent guarantee of provable integrity all the way from data acquisition until its de facto usage as an aggregated statistical result or global ML model. They also do not address the issues of input validation or state preservation for arbitrary functions $\sf{\mathcal{F}}$ to be executed on $\sf{\mathcal{P}rv}$.
@@
Private Data Collection on Edge Devices. Complementary to privacy mechanisms focusing on hiding private data from back-ends (e.g., LDP/FL), recent work has delved into assuring that private data is secure against compromised sensing devices “from its birth”, i.e., from the moment when it is digitized. VERSA [43] was proposed as a HW/SW architecture to guarantee that only the correct execution of expected and explicitly authorized software can access and manipulate sensing interfaces. As a consequence, it blocks malware/modified software from accessing sensitive sensed quantities by default. Following this notion, Sensing And Actuation As A Privilege (SA4P) [14, 15] realizes this concept using ARM TrustZone.